OpenStudy (anonymous):

-3t-4/p=6/s solve for P

OpenStudy (angela210793):

is it \[(-3t-4)\div p=6\div s?\]

OpenStudy (angela210793):

if it is then P=[s(-3t-4)]/6

OpenStudy (anonymous):

is like this : -3t-4/p= 6/s solve for "P"

OpenStudy (anonymous):

Ok .... \[L_AT^EX \:\:\:\:\:Magic\] \[\frac{-3t -4}{p} = \frac{6}{s}\] Tell me it must be like this ?

OpenStudy (anonymous):

Or

OpenStudy (anonymous):

\[-3t - \frac{4}{p} = \frac{6}{s}\]

OpenStudy (anonymous):

ninomeza answer please

OpenStudy (akshay_budhkar):

lol!! ishaan

OpenStudy (anonymous):

yes ishaan ..am new to this.. that correct typing on the problem

OpenStudy (anonymous):

Ishaan94 ... how do you solve it? please